OpenStudy (horsegirl27):

So, I did this homework problem but I got it wrong. I was wondering if someone could look at me work and explain it. Medal and fan :)

OpenStudy (horsegirl27):

http://prntscr.com/6uu07j here's the question

OpenStudy (horsegirl27):

2. w1*d1=w2*d2 w1*d1/w2 =w2/w2*d2 w1*d1/w2 =d2

OpenStudy (horsegirl27):

That's what I got, but my teacher said it was wrong and idk what to change.

OpenStudy (anonymous):

can you make g the subject of the formula?

OpenStudy (horsegirl27):

how?

OpenStudy (horsegirl27):

I'm confused how to do this problem, I wasn't sure so I just worked on it and did what I thought, but now I'm really confused :/

OpenStudy (anonymous):

You need to calculate g and you have to rearrange it in the formula in order to find it.

OpenStudy (horsegirl27):

okkkkkkkkk............

OpenStudy (anonymous):

\[g=\frac{ 2*d }{ t^2 }\] Got this?

OpenStudy (horsegirl27):

Sort of.

OpenStudy (horsegirl27):

oh wait, I think I posted the wrong work for the problem oops, let me give u what I actually had :P XD

OpenStudy (anonymous):

Right. The original formula is \[d=0.5g t^2\] which can also be written as \[d=\frac{ g t^2 }{ 2 }\]. Agree?

OpenStudy (horsegirl27):

ok, yeah that's what I had to start, do u want to see my real work lol?

OpenStudy (anonymous):

Yes, please.

OpenStudy (horsegirl27):

d= 0.5gt^2 d/0.5=0.5/0.5gt^2 d/0.5= gt^2 d/0.5t^2=g Is any of this correct?

OpenStudy (anonymous):

That is correct! Just replace the values and get the answer, right?

OpenStudy (anonymous):

What do you get @horsegirl27 ?

OpenStudy (horsegirl27):

oh, ok